OpenStudy (anonymous):

Find the domain and range of the function: h(x)= sqr root( 25 - x^2)

OpenStudy (anonymous):

domain: solve 25-x^2>0 so -5< x<5

OpenStudy (anonymous):

Range is the same: -5<y<5

OpenStudy (anonymous):

\[h(x)= \sqrt{( 25 - x^2)}\]\[D_h=\left\{ x \in \mathbb{R}/25-x^2\ge0 \right\}\]\[D_h=[-5,5]\]